# Large Language Models Are Cross-Lingual Knowledge-Free Reasoners

### v1

@@ Abstract
Large Language Models have demonstrated impressive reasoning capabilities across multiple languages. However, the relationship between capabilities in different languages is less explored. In this work, we decompose the process of reasoning tasks into two separated parts: knowledge retrieval and knowledge-free reasoning, and analyze the cross-lingual transferability of them. With adapted and constructed knowledge-free reasoning datasets, we show that the knowledge-free reasoning capability can be nearly perfectly transferred across various source-target language directions despite the secondary impact of resource in some specific target languages, while cross-lingual knowledge retrieval significantly hinders the transfer. Moreover, by analyzing the hidden states and feed-forward network neuron activation during the reasoning tasks, we show that higher similarity of hidden representations and larger overlap of activated neurons could explain the better cross-lingual transferability of knowledge-free reasoning than knowledge retrieval. Thus, we hypothesize that knowledge-free reasoning embeds in some language-shared mechanism, while knowledge is stored separately in different languages.  [^1] Our code and data is available at: https://github.com/NJUNLP/Knowledge-Free-Reasoning.
@@ 1 Introduction
Large language models (LLMs) today have shown strong multitask and multilingual performance in various domains Huang and Chang (2022), including robust reasoning capabilities across multiple languages Ranaldi and Zanzotto (2023), even for low-resource languages in the training corpus Shi et al. (2022).
@@
Previous study reveals that these multilingual LLMs possess certain ability of multilingual transfer Ye et al. (2023), which means the skills or knowledge learned with one language can be automatically transferred to another language without extra training. However, the effect of such cross-lingual transfer varies across tasks. In knowledge-related tasks, current LLMs show unsatisfactory cross-lingual transfer Gao et al. (2024), while in certain reasoning tasks, more effective transfer is observed Qi et al. (2023). Previous study lacks the analysis on the difference between these tasks, and does not dig further into the specific factors affecting the transfer effectiveness.
@@
In this study, we divide a reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning, and investigate the transfer effectiveness of them in current LLMs. The former means recalling information encountered during training, while the latter refers to organizing the given knowledge in the context into a final answer.
@@
Based on such definition, we perform evaluation and interpretability analysis on LLaMA-2 models Touvron et al. (2023) to compare the cross-lingual transferability of knowledge retrieval and knowledge-free reasoning. In the evaluation part, by controlling the requirement of knowledge retrieval in the StrategyQA (Geva et al., 2021) dataset and constructing a knowledge-free reasoning dataset, we compare the cross-lingual transfer effectiveness of knowledge-related and knowledge-free reasoning. In the interpretability analysis part, we assess the cross-lingual computational similarity of hidden states and Feed-Forward Network (FFN) neural activation to trace the computational process of cross-lingual transfer of knowledge-related and knowledge-free reasoning in LLMs. Our main findings are:
@@
  • Knowledge retrieval significantly hinders cross-lingual transfer. The more knowledge retrieval is required in the task, the lower effectiveness of cross-lingual transfer is observed.
@@
  • The knowledge-free reasoning ability can be effectively transferred to other languages after finetuning in one, while the model’s ability in these languages are less important.
@@
  • The overall cross-lingual computational similarity of knowledge-free reasoning is significantly higher than knowledge-related reasoning. Also, the difference in such similarity level is the most significant in the middle-high layers of the model. This suggests a language-shared mechanism of knowledge-free reasoning in multilingual LLMs.
2 Methods
The work in this paper consists of two parts: evaluation and interpretability analysis. For each part, we employ specific datasets and design certain metrics or measurements.
@@ 2.1 Datasets
This part introduces the datasets used in this paper for the analysis of the cross-lingual transfer of reasoning tasks. For datasets available only in English or part of the tested languages, we translate them into other languages with Google Translate.
2.1.1 Adapted StrategyQA dataset
The StrategyQA dataset is a multi-domain reasoning dataset including common sense, arithmetic, symbolic, and logical reasoning, where model is given at least two pieces of evidence from Wikipedia, and required to give "yes" or "no" answers to the questions. An example can be found in Table 3 in Appendix B.
Since Wikipedia is a commonly used source in LLM pretraining, the model can either collect the required information from evidence in the context, or retrieve it from its internal knowledge strorage. Thus, by masking out some of the evidence, we can control the need of knowledge retrieval in the reasoning process, which allows us to analyze the impact of knowledge retrieval on the cross-lingual transfer of reasoning skills. Namely, we design two kinds of scenarios in the experiments:
  • No Fact (NF): The model is given only the questions and must retrieve the necessary knowledge (evidence).
@@ 2.1.2 Knowledge-free reasoning dataset
Although the need for knowledge retrieval can be controlled with the adapted StrategyQA dataset, the questions themselves may still contain implicit requirement for knowledge retrieval. To rule out such effect and focus on the knowledge-free reasoning process, we construct a data that minimizes the need for knowledge retrieval. Inspired by Wei et al. (2022)’s taxonomy of reasoning tasks, we developed the Knowledge-Free Reasoning Dataset (KFRD), which consists of three fundamental reasoning tasks: arithmetic reasoning, symbolic reasoning, and logic reasoning in the 4-choice format. The dataset is constructed by filling GPT-4 Achiam et al. (2023) generated entities or simple English words into a unified template with one input part, one transformation rule, and one options part. Then, using the translated templates, entities and words, the dataset is expanded to all the tested languages. The template and examples can be found in Figure 11 and Table 1 in Appendix A.
Arithmetic Reasoning:
transforming one or two input numbers through mathematical operations into one or two output numbers. The mathematical operations include: basic arithmetic operations (addition, subtraction, multiplication, division), equality, geometric progression, arithmetic progression, and sorting.
Symbolic Reasoning:
transforming 3-5 words from the respective language input through symbolic operations to obtain the output. The symbolic operations include: repetition, addition, deletion, reordering, and their combinations, with up to three operations applied in combination.
Logical Reasoning:
selecting propositions derived from eight given premises using provided logical rules. The logical rules include: Implication Elimination, Conjunction Introduction, Conjunction Elimination, Disjunction Introduction, Disjunction Elimination, and Proof by Contradiction.
2.1.3 Knowledge-related reasoning dataset
To further compare the cross-lingual transfer of knowledge retrieval and knowledge-free reasoning, three knowledge-related reasoning datasets are used in the interpretability analysis. The three datasets are:
MKQA:
The MKQA dataset Longpre et al. (2021) is an open-domain question-answering evaluation set designed to provide a challenging benchmark for multilingual question-answering. The dataset consists of 10,000 question-answer pairs and covers 26 different languages.
BoolQ:
BoolQ (Clark et al., 2019) is a question answering dataset for yes/no questions, comprising 15,942 examples. The questions originate from queries to the Google search engine with Wikipedia pages as answers.
AmbigQA:
AmbigQA (Min et al., 2020) is an open-domain dataset for question answering that includes 14,042 questions. It addresses the issue of ambiguous answers by rephrasing the related questions.
Some examples of these datasets are shown in Table 3 in Appendix B.
@@ 2.2 Evaluation metric
In order to measure the model’s cross-lingual transferability while mitigating the difficulty difference in of the datasets, we calculate the Cross-lingual Transfer Ratio (XLTR) following the prior work Gao et al. (2024). The formula is as follows:
@@
where $s$ and $t$ denote the source and target languages in the transfer. $C_{x}$ represents the set of correct answers in language $x$, and $A_{r}$ is the accuracy of random choices for the given task. If after training on a task in one specific language (namely, English), the model shows an XLTR score close to 100% in other untrained languages, we say it achieves Fully Cross-Lingual Transfer (FCLT) in this task.
@@ 2.3 Interpretability measurements
We employ two metrics to evaluate the cross-lingual computational consistency of the LLM during knowledge-related and knowledge-free reasoning processes.
@@ 2.3.2 Neuron Activation Overlap (NAO)
The Neuron Activation Overlap refers to the extent of overlap between activated neurons across multiple languages for the same corpus. The calculation steps are as follow:
  1. 1.
    Input a question in multiple languages into the model and extract the activation values of the last token in the question in each language.
  2. 2.
    Identify activated neurons. A neuron is considered activated if the absolute value of its activation exceeds a certain threshold.
  3. 3.
    Calculate the overlap of activated neurons in all the tested languages.
For a question sample $x$, the formula is as follows:
@@
3 Experiment Settings
3.1 Language and model choice
Language choice
To capture a broad spectrum of linguistic diversity, we selected ten languages based on their extensive use and representation of diverse linguistic families, following Gao et al. (2024). The languages selected are English (en), German (de), French (fr), Italian (it), Russian (ru), Polish (pl), Arabic (ar), Hebrew (he), Chinese (zh), and Japanese (ja). Further details are provided in Appendix C.
Model choice
We selected several LLMs including LLaMA-2-7B-Chat (Touvron et al., 2023), BLOOMZ-MT-7B (Muennighoff et al., 2023), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023), and Qwen-1.5-7B-Chat (Bai et al., 2023) to conduct transferability experiments on existing and synthetic datasets.
For experiments focusing on the impact of the training language (Section 4.3.1) and interpretability (Section 4.4), we used LLaMA-2-7B-Chat as a representative model for analysis.
For the experiments evaluating impact of the target language proficiency in specific directions (Section 4.3.2), we selected derived models of LLaMA-2-7B and Mistral-7B to conduct experiments on Arabic and Hebrew respectively, both of which are low-resource languages.
3.2 Fine-tuning settings
We perform LoRA fine-tuning (Hu et al., 2021) on all model blocks in all experiments due to the limited computational resources. More details about fine-tuning can be found in Appendix E.
3.3 Decoding settings
In all experiments, we perform constrained decoding to prevent the model from generating tokens other than the desired choices (Yes/No for StrategyQA, A/B/C/D for KFRD).
To ensure consistency in the final token across different datasets, we added a language-specific ’?’ to certain datasets. For details on these modifications and metrics, please refer to Appendix D.
4 Results
Figure 1: XLTR of different models on StrategyQA. Solid lines: WF results; Dashed lines: NF results. The label of training language (en) is capitalized.
Figure 2: XLTR of LLaMA-2-7B-Chat on StrategyQA under different settings. NF stands for the No Facts setting, while WF-1 and WF-2 indicate that one or two facts are given for each question, respectively. WF-all signifies that all facts are provided.
Figure 3: XLTR on the different parts of KFRD
Figure 4: Accuracy of LLaMA-2-7B-Chat from different languages on three parts of KFRD
Figure 5: XLTR of LLaMA-2-7B-Chat from different languages on three parts of KFRD
4.1 Impact of knowledge retrieval on cross-lingual transfer
We analyze the impact of the amount of retrieved knowledge on cross-lingual transfer in different settings of the StrategyQA dataset. The results for cross-lingual transfer ratio are presented in Figure 1, while the accuracy results are detailed in Figure 12 in Appendix G.
Knowledge retrieval requirement harms cross-lingual transfer
The experimental results indicate that, for all languages, the cross-lingual transfer ratio of all models are significantly higher when the necessary knowledge for reasoning is provided compared to when it is not. This suggests that the requirement for knowledge retrieval significantly hinders the model’s cross-lingual transferability when solving reasoning problems.
Less knowledge retrieval makes better cross-lingual transfer
Additionally, we conduct a more detailed evaluation using the LLaMA-2-7B-Chat model to observe the changes in cross-lingual transfer ratio with varying amounts of knowledge entries. As shown in Figure 2, the experimental results demonstrate that the transfer rate increases as the number of provided knowledge entries increases from none to all required knowledge. This further validates the conclusion that the retrieval of more knowledge significantly impacts cross-lingual transferability.
4.2 The cross-lingual transfer of knowledge-free reasoning
We assess the cross-lingual transferability of the model’s knowledge-free reasoning capabilities. We evaluate the performance of the KFRD by training in English and transferring to the other 9 languages. The resulting cross-lingual transfer ratio are shown in Figure 3, and accuracy results are shown in Figure 13 in Appendix G.
The results demonstrate that the KFRD exhibits extremely high cross-lingual transfer performance for most language pairs. For 7 out of the 9 languages, it can be observed that the cross-lingual transfer performance in knowledge-free reasoning tasks often exceeds 90%, with some instances approaching 100%, thereby achieving near FCLT.
However, for the two lowest-resource languages Arabic and Hebrew, the cross-lingual transferability is significantly poorer. We hypothesize that this may be due to the model’s weaker language proficiency in these two languages, which adversely affects its transferability. Further analysis on this issue is provided in the next section.
4.3 Impact of language proficiency on cross-lingual transfer
In this section, we explore the impact of language proficiency of the training and target language with additional experiments.
4.3.1 Training language proficiency
To test the impact of training language proficiency, we select German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages as training languages. Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages. The accuracy results are shown in Figure 5, and the transfer ratio is presented in Figure 5.
The results show that, though training with low-resource languages resulted in slightly lower accuracy (which is particularly noticeable in the arithmetic dataset), the models show no significant differences in transfer ratio when trained with high-resource or low-resource languages, indicating that the proficiency and resource of the training language has no significant effect on the cross-lingual transfer of knowledge-free reasoning.
4.3.2 Target language proficiency in specific directions
In previous experiments, we observe strong cross-linguistic transfer performance between most languages. However, the transferability from English to Arabic and Hebrew was significantly weaker. We hypothesize that this is related to the model’s weaker language proficiency in these two target languages.
In this section, we select models from Hugging Face that have undergone Continuous Pre-Training (CPT), Supervised Fine-Tuning (SFT), and a combination of both (CPT + SFT) on the LLaMA-2 or Mistral platforms. These adapted models have better proficiency in the respective languages. The selected models are listed in Table 2 in Appendix F.
We train a Vanilla model and the above fine-tuned models on the KFRD in English and test their transferability from English to the respective target languages. The transfer rate results are shown in Figure 6, and the accuracy results are provided in Figure 14 in Appendix G.
The experimental results indicate that the Vanilla model exhibits very low transfer rates for low-resource languages. However, after applying CPT, SFT, or CPT+SFT, the transfer rates increase significantly. Notably, for Hebrew, the transfer rate reached over 95%, approaching FCLT. This suggests that the proficiency in Arabic and Hebrew limits the cross-linguistic transfer of knowledge-free reasoning abilities from English to these languages. Improving proficiency in the target language can alleviate this limitation. This demonstrates that the cross-linguistic transfer of knowledge-free reasoning in specific directions is influenced by target language proficiency.
4.4 Analyzing differences in transfer characteristics

### v2

@@ Abstract
Large Language Models have demonstrated impressive reasoning capabilities across multiple languages. However, the relationship between capabilities in different languages is less explored. In this work, we decompose the process of reasoning tasks into two separated components: knowledge retrieval and knowledge-free reasoning, and analyze the relationship between cross-lingual transferability and these two components. With adapted commonsense reasoning datasets and constructed knowledge-free reasoning datasets, we show that the knowledge-free reasoning capability can be nearly perfectly transferred across various source-target language directions despite the secondary impact of resource in some specific target languages, while cross-lingual knowledge retrieval significantly hinders the transfer. Moreover, by analyzing the hidden states and feed-forward network neuron activation during the reasoning, we show that higher similarity of hidden representations and larger overlap of activated neurons could explain the better cross-lingual transferability of knowledge-free reasoning than knowledge retrieval. Thus, we hypothesize that knowledge-free reasoning shares similar neurons in different languages for reasoning, while knowledge is stored separately in different languages.
 [^1] Our code and data is available at: https://github.com/NJUNLP/Knowledge-Free-Reasoning.
@@ 1 Introduction
Large language models (LLMs) today have shown strong multitask and multilingual performance in various domains Huang and Chang (2022), including robust reasoning capabilities across multiple languages Ranaldi et al. (2024), even for low-resource languages in the training corpus Shi et al. (2022).
@@
Previous study reveals that these multilingual LLMs possess certain ability of multilingual transfer Qi et al. (2023); Gao et al. (2024); Ye et al. (2023), which means the skills or knowledge learned with one language can be automatically transferred to another language without extra training. However, the effect of such cross-lingual transfer varies across tasks. In certain tasks, especially knowledge retrieval tasks, current LLMs show unsatisfactory cross-lingual transfer Qi et al. (2023); Gao et al. (2024), while in certain reasoning tasks, more effective transfer is observed Ye et al. (2023). Previous study lacks the analysis on the difference between these tasks, and does not dig further into the specific factors affecting the transfer effectiveness.
Figure 1: Cross-lingual transfer involves training a model in one language and evaluating it in another. In this figure, the scenario depicts training in English. Reasoning tasks encompass both knowledge retrieval and knowledge-free reasoning. The cross-lingual transfer ratio is significantly lower for knowledge retrieval tasks (e.g., ZH case in EN: "Crocodiles, alligators, and pigeons are dangerous animals") compared to knowledge-free reasoning tasks, which transfer well across languages (e.g., ZH case in EN: "22 plus 23 equals 45").
@@
In this study, we divide a general reasoning task into two separated components: knowledge retrieval and knowledge-free reasoning. The former means recalling certain information from pre-training, while the latter refers to organizing the given knowledge in the context to perform inference and derive a final answer. Figure 1 provides a clearer understanding of these two components and illustrates the cross-lingual transfer issues explored in this paper.
@@
This paper includes both an evaluation part and an interpretability analysis part. In the evaluation part, we focus on the impact of knowledge retrieval component on cross-lingual transfer in reasoning tasks, and the transferability of knowledge-free reasoning capability, by adapting existing reasoning datasets with different levels of knowledge retrieval demand and creating a clean knowledge-free reasoning dataset, which only includes the knowledge-free reasoning component. In the interpretability analysis part, we assess the cross-lingual computational similarity of hidden states and Feed-Forward Network neuron activation to trace and compare the computational process of cross-lingual transfer of knowledge retrieval and knowledge-free reasoning components in LLMs. Our main findings are:
@@
  • Retrieval component significantly hinders cross-lingual transfer of reasoning tasks. The more knowledge retrieval is required in the task, the lower effectiveness of cross-lingual transfer is observed.
@@
  • The ability of knowledge-free reasoning component can be near-perfectly transferred to other languages after fine-tuning in one, while the model’s language proficiency in the target languages is also important.
@@
  • The overall cross-lingual computational similarity for knowledge-free reasoning tasks is significantly higher than for knowledge retrieval tasks, especially in the middle-high layers, which are primarily used for reasoning (Zhao et al., 2024; Wendler et al., 2024). This suggests a language-shared reasoning mechanism in multilingual LLMs.
2 Evaluation Methodology
2.1 Overview
Our evaluation focuses on two main aspects:
Impact of Knowledge Retrieval Demand on Cross-Lingual Transfer in Reasoning Tasks
We aim to analyze how varying levels of knowledge retrieval demand affect cross-lingual transfer in reasoning tasks. For this purpose, we leverage the commonsense reasoning datasets that provide questions along with several facts required to answer them. By controlling the number of facts provided to the model within the context, we can manipulate the levels of demand for knowledge retrieval. As more facts are provided, the model relies less on its internal knowledge storage. This controlled setup enables us to analyze how the demand for knowledge retrieval influences the cross-lingual transfer of the overall reasoning abilities.
Cross-Lingual Transfer of Knowledge-Free Reasoning
We also aim to investigate the cross-lingual transfer of knowledge-free reasoning, which is less explored in previous work. However, existing reasoning datasets often contain some degree of knowledge retrieval. For instance, while StrategyQA provides knowledge required for reasoning, it is not always complete. Similarly, certain mathematical datasets, like ASDiv, require knowledge retrieval for some problems (as demonstrated in Appendix G). This introduces noise when evaluating the cross-lingual transfer of knowledge-free reasoning. To address this, we constructed a new dataset, the Knowledge-Free Reasoning Dataset (KFRD), which entirely eliminates the need for knowledge retrieval. In addition, we selected several existing datasets that, to the best extent possible, meet the requirements of knowledge-free reasoning to further validate our conclusions. A more detailed explanation of why we constructed KFRD and the dataset selection criteria can be found in Appendix G.
@@ 2.2 Datasets
This section introduces the datasets used for evaluation. More details on the datasets and the construction process are in Appendix A.
2.2.1 Reasoning dataset with variable knowledge retrieval demand
We adapt the popular commonsense reasoning datasets, StrategyQA (Geva et al., 2021) and QASC (Khot et al., 2020), to analyze the impact of knowledge retrieval on cross-lingual transfer. They provide pieces of evidence from Wikipedia for answering the question. Examples can be found in Table A5.
Namely, we design two kinds of scenarios with variable knowledge retrieval demand in the experiments:
  • No Fact (NF): The model is given only the questions.
@@ 2.2.2 Knowledge-free reasoning dataset
Inspired by Wei et al. (2022)’s taxonomy of reasoning tasks, we developed the KFRD, which consists of three fundamental reasoning tasks: arithmetic reasoning (e.g., addition, subtraction, and other mathematical operations), symbolic reasoning(e.g., deletion, reordering, and other symbolic operations), and logic reasoning(e.g., Implication Elimination and other basic logical rules) . It is designed to evaluate a broad spectrum of knowledge-free reasoning and cross-lingual transfer performance.
We utilized GPT-4 Achiam et al. (2023) to generate multilingual parallel templates and fictitious entities, followed by manual verification. We then used code to generate the dataset based on these templates, entities, and predefined rules. This approach ensures that the tasks can be completed without requiring additional knowledge and guarantees the correctness of the QA pairs. The templates are multiple-choice questions, each composed of one input part, one transformation rule, and one options part. The examples and template are provided in Table 1 and Figure A1.
We also use the ASDiv Miao et al. (2021), Coin Flip Wei et al. (2022), and ProofWriter Tafjord et al. (2020) dataset as a representation of arithmetic, symbolic, and logical reasoning to further validate our conclusions.
Arithmetic Reasoning
Input 11, 645 (two numbers)
Transformation Rule Addition (a mathematical operation)
Output Options A) 595  B) 536 C) 771  D) 656
Symbolic Reasoning
Transformation Rule Swap the positions of the 5th and 2nd words; Delete the 2nd word (1-3 symbolic operations)
Logical Reasoning
Input Alex is Aurora Vale. Everything that is Aurora Vale is Omicron Delta. Stella is not Chronos Wasteland. Max is not Dreamweaver’s Haven. Suppose Sally is Whispering Meadows, then Sally is Chimerical Citadel. Everything that is Ebonwyrm Abyss is Phoenixfire Ridge. (6 propositions)
Transformation Rule Implication Elimination (a logical rule)
Table 1: Examples of different tasks in the KFRD dataset
@@ 2.3 Evaluation metric
In order to assess the model’s cross-lingual transferability, we select the Cross-lingual Transfer Ratio (XLTR) as the evaluation metric, following Gao et al. (2024). The formula is as follows:
@@
where $s$ and $t$ denote the source and target languages in the transfer. $C_{x}$ represents the set of correct answers in language $x$, and $A_{r}$ is the accuracy of random choices for the given task.
If the model shows an XLTR score close to 100% in a language direction, we say it achieves fully cross-lingual transfer in this direction.
We also evaluate the accuracy of models before fine-tuning on all datasets and find poor performance, suggesting that most of the model’s ability on transferred languages stem from cross-lingual transfer.
3 Experiment Settings
3.1 Language and model choice
Language choice
To capture linguistic diversity, we selected ten languages based on their extensive use and representation of diverse linguistic families, following Gao et al. (2024). The languages selected are English (en), German (de), French (fr), Italian (it), Russian (ru), Polish (pl), Arabic (ar), Hebrew (he), Chinese (zh), and Japanese (ja). Unless specified, we fine-tune the model in English and evaluate it on other languages. Further details are provided in Appendix B.
Model choice
We selected several LLMs, including LLaMA-2-7B-Chat (Touvron et al., 2023), BLOOMZ-MT-7B (Muennighoff et al., 2023), Mistral-7B-Instruct-v0.1 (Jiang et al., 2023), and Qwen-1.5-7B-Chat (Bai et al., 2023), for our experiments. To optimize resource use and demonstrate results clearly, we used LLaMA-2-7B-Chat (Touvron et al., 2023) as a representative model for some analyses.
3.2 Fine-tuning and decoding settings
We perform LoRA fine-tuning (Hu et al., 2021) on all model blocks in all experiments due to the limited computational resources. More details about fine-tuning can be found in Appendix D.
For decoding, we use constrained decoding in all experiments to ensure the model generates only the desired options (e.g., Yes/No for StrategyQA, A/B/C/D for KFRD).
4 Results
Figure 2: XLTR of different models on StrategyQA. Solid lines: WF-all results; Dashed lines: NF results. The label of training language (en) is capitalized.
Figure 3: XLTR of LLaMA-2-7B-Chat on StrategyQA under different settings.
Figure 4: XLTR on the different parts of KFRD
Figure 5: XLTR on the existing pseudo knowledge-free reasoning datasets
Figure 6: XLTR of LLaMA-2-7B-Chat on three parts of KFRD. The different lines indicate different trained languages.
4.1 Impact of knowledge retrieval demand on cross-lingual transfer
We analyze the impact of the amount of knowledge retrieve on cross-lingual transfer in different settings of the reasoning dataset. The results of StrategyQA for the cross-lingual transfer ratio are presented in Figure 2, while the accuracy results are detailed in Figure A2.
Knowledge retrieval requirement harms cross-lingual transfer
The experimental results indicate that, for all languages, the cross-lingual transfer ratio of all models are significantly higher when the necessary knowledge for reasoning is provided compared to when it is not. This suggests that the requirement for knowledge retrieval significantly hinders the model’s cross-lingual transferability when solving reasoning tasks.
More knowledge retrieval lowers cross-lingual transfer
We further conduct detailed evaluations using the LLaMA-2-7B-Chat model to observe the changes in cross-lingual transfer ratios as the amount of knowledge provided varies. As shown in Figure 3, the experimental results demonstrate that the transfer ratio decreases as the demand for knowledge retrieval increases. This further validates the conclusion that the retrieval of more knowledge significantly impacts cross-lingual transferability.
The results on the QASC dataset were consistent with those mentioned above. Detailed results can be found in Figure A3 and A4.
4.2 The cross-lingual transfer of knowledge-free reasoning
We assess the cross-lingual transferability of the model’s knowledge-free reasoning capabilities by evaluating the performance on KFRD and three corresponding existing reasoning datasets. The resulting cross-lingual transfer ratios are shown in Figures 4 and 5, while the accuracy results are presented in Figures A5 and A5.
The results demonstrate that the KFRD exhibits extremely high cross-lingual transfer performance for most language pairs. For 7 out of the 9 languages, it can be observed that the cross-lingual transfer ratio in knowledge-free reasoning tasks often exceeds 90%, with some instances approaching 100%, thus achieving near-full cross-lingual transfer. Moreover, results from three existing datasets further validate this finding, showing that all models achieve satisfactory transfer ratios across high-resource languages.
For some low-resource languages, such as Hebrew and Arabic in LLaMA-2, German and Hebrew in BLOOMZ  [^2] For the language distribution, please see Appendix F., the cross-lingual transferability is significantly poorer. We hypothesize that this may be due to the model’s weaker language proficiency in these languages, which negatively impacts its transferability. Further analysis of this issue is provided in the following section.
It is noticeable that there are still minor differences in XLTR between KFRD and the existing datasets in the arithmetic reasoning and logical reasoning tasks. However, these differences do not affect the overall conclusion.
We manually check the data samples and find that there are some interfering cases that can affect cross-lingual transfer, such as minor knowledge retrieves, translation issues, and counterfactual information, as detailed discussed in the Appendix G.
4.3 Impact of language proficiency on cross-lingual transfer
4.3.1 Training language proficiency
To evaluate the impact of training language proficiency, based on the data distribution of LLaMA-2 (see Appendix F) and previous experiments, we selected German and Chinese as representatives of high-resource languages, and Arabic and Hebrew as representatives of low-resource languages for training. Then, we train models on the KFRD in these languages and evaluated their performance across the 10 languages. The resulting cross-lingual transfer ratios are presented in Figure 6, while the accuracy results are shown in Figure A6.
The results show that the models show no significant differences in transfer ratio when trained with high-resource or low-resource languages, indicating that the proficiency and resource of the training language has no significant effect on the cross-lingual transfer of knowledge-free reasoning.
4.3.2 Target language proficiency
In previous experiments, we observe the transferability from English to Arabic and Hebrew was significantly weaker in LLaMA-2 and Mistral. We hypothesize that this is related to the model’s weaker language proficiency in these two target languages.
In this section, we select models from Hugging Face that have undergone Continual Pre-Training (CPT), Supervised Fine-Tuning (SFT), and a combination of both (CPT + SFT) on the LLaMA-2 or Mistral platforms. These adapted models have better proficiency in the respective languages. The selected models are listed in Table A1.
The transfer ratio results of the vanilla and the above fine-tune models are shown in Figure 7, and the accuracy results are provided in Figure A7.
Figure 7: Averaged XLTR from English to Arabic/Hebrew across three parts of our KFRD dataset for models in different stages trained in Arabic/Hebrew
The results indicate that the vanilla model exhibits very low transfer rates for low-resource languages. However, after applying CPT, SFT, or CPT+SFT, the transfer ratio increases significantly. Notably, for Hebrew, the transfer ratio reach over 95%, achieving fully cross-lingual transfer. This suggests that proficiency in Arabic and Hebrew limits the cross-linguistic transfer of the knowledge-free reasoning component, while improving proficiency in the target language can alleviate this limitation.
5 Interpretability Analysis
5.1 Overview
Built on previous research (Hu et al., 2024; Gao et al., 2024) and our experiments, we observed that the cross-lingual transferability of knowledge retrieval ability is significantly weaker than that of knowledge-free reasoning. To better understand the reasons behind this difference, we conducted a detailed analysis on model internals using two methods: Cosine Similarity of Hidden States and Neuron Activation. Both of the methods have been widely used to measure text similarity (Reimers and Gurevych, 2019; Malkiel et al., 2022; Wang et al., 2024) and analyze models (Dalvi et al., 2019; Sajjad et al., 2022; Rai and Yao, 2024).
@@ 5.2 Interpretability measurements
This section introduces the measurements used for interpretability analysis. Further details for these metrics are in Appendix C.
@@ 5.2.2 Neuron Activation Overlap (NAO)
Neuron Activation Overlap measures the extent of shared neuron activations across languages for the same input.
To calculate NAO, we input a question in multiple languages, extract the neuron activation values of the last token of the input, and identify the neurons whose absolute values surpass a set threshold, labeling them as activated. Then their overlap (NAO) is computed as follows for a question sample $x$:
@@
5.3 Knowledge retrieval dataset
We selected MKQA Longpre et al. (2021), BoolQ (Clark et al., 2019), and AmbigQA (Min et al., 2020) as representative datasets of knowledge retrieval tasks for the interpretability analysis. Most questions in these datasets can be answered through a single instance of knowledge retrieval. Examples of these datasets are shown in Table A8.
5.4 Interpretability results